no fice s as the cortte cools esS cun thus ddecreses cannot be CEee ls in a 5,00 c aping up nlorel is achally lue h so ke of 33 °c in avery well-nslate
Answer #1

Solution 1 (b)

Let us go to the basics first.

We shall use the concept of heat balance.

Heat absorbed by ice to reach from -30 to 0 degree celsius

= mcdt (c = 2108 J/kg degree celsius = specific heat of ice)

= 0.055 * 2108 * (0 - ( - 30))

= 3478.2 J

Now, heat absorbed by ice at 0 degree celsius to melt at 0 degree celsius

= m* L (L= latent heat of fusion = 3.33 x 105 J/kg)

= 0.055 kg * 3.33 x 105 J/kg

= 18315 J

Now let the final temperature be T.

So, total heat absorbed by ice = heat released by water

Thus, 3478.2 + 18315 + 0.055 * 4187 * (T - 0)

= 0.45 * 4187 * (35 - T)

Solving this equation gives

T = 20.88 degree celsius (Answer == Final temperature)

Thanks!!!